OpenStudy (anonymous):

Please help on this problem What is (48 + 52)² = 48² + 2(48)(52) + 52²?

OpenStudy (jiteshmeghwal9):

\(\Huge{(a+b)^2=a^2+b^2+2ab}\)

OpenStudy (anonymous):

(48 + 52)² = 48² + 2(48)(52) + 52² is simply (a+b)² + 2ab + b², where a = 48, b =52 If you want to get the value, you can simply add 48+52, then take the square of the answer. It is a nice answer!

OpenStudy (jiteshmeghwal9):

this question use this identity :)

OpenStudy (anonymous):

so is the answer 5144

OpenStudy (anonymous):

What is 48+52?

OpenStudy (anonymous):

100

OpenStudy (anonymous):

what is 100²?

OpenStudy (jiteshmeghwal9):



OpenStudy (jiteshmeghwal9):

\(100^2=100\times100\)

OpenStudy (jiteshmeghwal9):

what is 100*100 ?

OpenStudy (jiteshmeghwal9):

@togoodtobetrue4u ??

OpenStudy (anonymous):

100,000 right

OpenStudy (anonymous):

Not really.

OpenStudy (anonymous):

Then I'm confused

OpenStudy (anonymous):

100 x 100 = ...?!

OpenStudy (anonymous):

sorry it is 10,000

OpenStudy (anonymous):

Yes :)